OpenStudy (anonymous):

Question: What is the system of substitution for the following problem; 3x-y=11 & 5y-7x=1

OpenStudy (anonymous):

I mean by or of substitution!

OpenStudy (texaschic101):

3x - y = 11 3x - 11 = y now sub 3x - 11 in for y in the other equation

OpenStudy (texaschic101):

are you supposed to solve it ?

OpenStudy (anonymous):

Yes, I am. I just got 8x=56 and then you would get x=7! Then you would plug &chug the 7 into the 3x-11=y!

OpenStudy (anonymous):

Then 11=y and then the ordered pair answer is (7,11) and then you should plug and chug again to check if you are right! ( into the pre-existing equations I had trouble with!)

OpenStudy (texaschic101):

5y - 7x = 1 5(3x - 11) - 7x = 1 15x - 55 - 7x = 1 15x - 7x = 1 + 55 8x = 56 x = 7 3x - 11 = y 3(7) - 11 = y 21 - 11 = y 10 = y check.. 5y - 7x = 1 5(10) - 7(7) = 1 50 - 49 = 1 1 = 1 (correct) answer is : x = 7 and y = 10 you are correct :)

OpenStudy (anonymous):

Thank you for helping me with Algebra, I do award you!

OpenStudy (texaschic101):

anytime :)